# Temperature Dependence of Heavy Quark Diffusion from (2+1)-flavor Lattice QCD

### v1

@@ 1 Introduction
The study of heavy quarks in the quark-gluon plasma (QGP) [1, 2, 3] is essential for understanding the general properties of strongly interacting matter produced in heavy-ion collisions (HIC). Heavy quarks, such as charm and bottom quarks, provide unique insights into the non-equilibrium dynamics of the QGP, as they are produced in the early stages of the collision and then propagate throughout the evolving plasma. Therefore, the study of interaction and equilibration of heavy quarks is crucial for understanding many properties of the QGP, such as its near-perfect fluid behavior [4, 5, 6].
@@
The interaction of heavy quarks with the medium can be effectively described by Langevin dynamics [7] for sufficiently large masses, and their motion is characterized with two transport coefficients: the heavy-quark momentum diffusion coefficient $\kappa$ and the drag coefficient $\eta$. These transport coefficients play a central role, since they determine the momentum distributions of the heavy-flavor hadrons in the HIC, which have been precisely measured at RHIC [8, 9] and LHC [10].
@@
The momentum diffusion coefficient $\kappa$ quantifies the average momentum transferred from the QGP to a heavy quark by random, uncorrelated in time, momentum kicks, with $3\kappa$ being the mean squared momentum transfer per unit time. Physically, it describes how much momentum a heavy quark loses and how fast it thermalizes in the plasma. For sufficiently heavy quarks, the relaxation time is expected to be $M/T$ times larger than that of light quarks [11, 7, 12], where $M$ is the mass of the heavy quark and $T$ is the temperature of the medium. Recent measurements have confirmed that this relaxation time is indeed shorter than the mean lifetime of the QGP (even with the enhancement of the prefactor $M/T$), suggesting a rather fast relaxation of heavy quarks and confirming the strongly interacting nature of the matter produced in HIC. Therefore, because of its relevance to the phenomenology of heavy ions, an accurate estimate of $\kappa$ has become a critical goal for both theoretical and experimental studies.
@@
Lattice field theory techniques provide a first-principles approach to the non-perturbative aspects of quantum chromodynamics (QCD), including the extraction of $\kappa$. However, the calculation is far from trivial, since it requires a spectral reconstruction of the correlation functions computed numerically on the lattice. Like all transport coefficients, the diffusion coefficient is directly encoded in the low-frequency behavior of the spectral function. It is precisely this inversion problem that makes the extraction extremely challenging, i.e. even when the correlators are determined precisely, the spectral functions extracted from them —and consequently the diffusion coefficient— can still carry large uncertainties.
@@
The heavy-quark diffusion coefficient has been calculated in perturbation theory [11, 7, 13], but the calculation is reliable only for asymptotically large temperatures $T$. In the present work, we build on previous studies to extract the heavy-quark diffusion coefficient from lattice QCD [14, 15, 16, 17, 18, 19, 20, 21, 22]. The HotQCD collaboration calculated $\kappa$ in a (2+1)-flavor lattice QCD framework for physical strange quarks and rather unphysical light quarks with mass $m_{l}=m_{s}/5$ [21, 20] corresponding to the pion mass of $m_{\pi}\simeq 320$ MeV, and for a temperature range between $T=195$ and $352$ MeV. The effect of light-quark masses is expected to be irrelevant for sufficiently large temperatures, but important for temperatures below $250$ MeV. Thus, the main goal of this work is to extend the previous calculation to an almost physical point, with light-quark masses of $m_{l}=m_{s}/20$ corresponding to a pion mass of $m_{\pi}\simeq 160$ MeV, and for a wider temperature range, from values close to the crossover temperature $T_{c}$ up to the GeV scale. The extension to high temperature is needed for making contact with perturbation theory.
@@ 2 Theoretical framework
Close to equilibrium, the response of strongly interacting matter to small perturbations can be studied in terms of a handful of transport coefficients, such as the charge diffusion coefficient, or the shear and bulk viscosities. These coefficients can be extracted from current-current correlation functions, where the type of current determines the specific coefficient to be computed. Among all the coefficients, the easiest one to extract is the heavy-quark momentum diffusion coefficient $\kappa$, which is formally defined as the mean-squared momentum transfer per unit time experienced by a heavy quark immersed in a QGP. Together with the equilibrium mean-squared momentum that the heavy quark carries in the medium, this determines how long it takes for the quark to thermalize in the QGP. For a more extended review see section 6 of [23].
@@
$G^{ii}(\tau,\mathbf{p},T)=\int\mathrm{d}\mathbf{x}\exp{i\mathbf{x}\mathbf{p}}\langle J^{i}_{V}(0,0,T)J^{i}_{V}(\tau,\mathbf{x},T)\rangle=\int_{0}^{\infty}\frac{\mathrm{d}\omega}{\pi}\rho^{ii}(\omega,\mathbf{p},T)K(\omega,\tau,T)\,,$ (2.1)
@@
where $\chi_{q}$ is the heavy quark number susceptibility, $q=c,b$ for charm and bottom, respectively. The determination of $D_{s}$ on the lattice through Eq. (2.3) is extremely challenging, as discussed in Ref. [24]. An alternative approach relies on the use of the heavy-quark effective theory (HQET) [25] as discussed in Refs. [26, 27, 28]. In the HQET approach, the current-current correlation function is expressed in terms of the chromo-electric and chromo-magnetic correlation functions [26, 27]:
@@
The $\omega\rightarrow 0$ limit of the corresponding spectral functions defines the transport coefficients $\kappa_{E}$ and $\kappa_{B}$ [26, 27]:
@@
Thus, the heavy-quark momentum diffusion coefficient, up to corrections ${\cal O}(T^{2}/M^{2})$, can be written as [27]:
@@
The chromo-electric and chromo-magnetic correlation functions are constructed in terms of gluonic operators, which suffer from strong short-distance fluctuations that yield low signal-to-noise ratio. Accurate estimation of the correlation functions (2.4) and (2.5) thus requires noise reduction techniques, such as multi-level updating [29]. While the electric correlation function requires only a finite renormalization at finite lattice spacing, and is regulator independent within the ${\overline{\rm MS}}$ scheme [30], the magnetic correlator involves a non-trivial anomalous dimension [25], and therefore it requires a specific renormalization method. On the lattice, we will use the gradient flow (GF) renormalization procedure [31, 32, 33, 34] to solve both problems at once, which has been shown to be as good as multi-level algorithms for noise reduction in the quenched theory [16, 18], while providing renormalized field operators for any flow time $\tau_{F}$ larger than the square of the lattice-spacing scale. We will explain our choice of renormalization method as well as a small description of our GF implementation in Sec. 3.
In both, the electric and the magnetic case, spectral reconstruction is a crucial and difficult aspect of the calculation. The main problem arises from the fact that the correlation functions are computed only at some discrete Euclidean time intervals $\tau$. Thus, to extract the spectral function $\rho(\omega,T)$ from these data, one has to use a theoretically motivated Ansatz. For this purpose we can use the perturbative QCD (pQCD) estimates of the spectral function. At LO and NLO [26], the spectral function in pQCD is predicted to behave as $\rho(\omega,T)\propto\omega^{3}$ for large $\omega$, while the low-energy behavior of the spectral function directly yields the diffusion coefficient $\kappa$ as $\rho(\omega,T)\propto\kappa_{E,B}\omega$. To interpolate between these two regions, we will explore several methods, like for instance sharp, smooth, or polynomial transition between the IR and UV behavior of the spectral function. These methods are discussed in detail in Sec. 5. In order to determine $\kappa$ and $D_{s}$, we also need to determine the average thermal velocity, ${\langle\mathbf{v}^{2}\rangle}$ and averaged thermal momentum, ${\langle\mathbf{p}^{2}\rangle}$ of the heavy quark. Following previous works [21, 20], in Sec. 5.3.2 we will determine these quantities in terms of a quasi-particle model (QPM).
@@ 3.1 Lattice Setup
$T$ [MeV] $\beta$ $a$ [fm] $a\crossproduct m_{s}$ $a\crossproduct m_{l}$ $N_{\sigma}$ $N_{\tau}$ conf. streams
@@
$T$ [MeV] $\beta$ $a$ [fm] $a\crossproduct m_{s}$ $a\crossproduct m_{l}$ $N_{\sigma}$ $N_{\tau}$ conf. streams
@@
Table 2: Input parameters of the lattice ensembles corresponding to light-quark masses $m_{l}=m_{s}/5$. We show in every case the temperature in MeV, the bare coupling with its corresponding lattice spacing in fm, the geometry $N_{\sigma}^{3}\crossproduct N_{\tau}$ and the number of configurations separated by 10 MC trajectories. The highlighted pink regions correspond to ensembles generated in previous works of the HotQCD collaboration [20, 21] or the TUMQCD collaboration [35].
We have performed lattice calculations in (2+1)-flavor QCD with physical strange-quark mass $m_{s}$ over a temperature range from $T=149$ MeV to $10$ GeV using the Highly Improved Staggered Quarks (HISQ) action [36] for the fermion sector and the tree-level improved Lüscher-Weisz action [37, 38] for the gauge fields. The calculations have been performed on $N_{\sigma}^{3}\crossproduct N_{\tau}$ boxes for two values of the light-quark mass: $m_{l}=m_{s}/20$, which corresponds to the pion mass of $160$ MeV in the continuum limit (and thus is close to the physical point), and $m_{l}=m_{s}/5$, corresponding to a pion mass of $320$ MeV. The lattice strange-quark mass was obtained from the parameterization of the line of constant physics (LCP) of [39]. The lattice spacing, and thus the temperature scale, has been fixed through the $r_{1}$ scale determined in Refs. [39, 40] with the value $r_{1}=0.3106(18)$ fm [41]. Though a very recent determination of $r_{1}$ in physical units yielded a smaller central value [42], it is still compatible with the scale used in this paper.
The gauge configurations were generated using the usual Rational Hybrid Monte Carlo (RHMC) algorithm [43, 44], where each configuration is saved after 10 Monte Carlo (MC) trajectories with an acceptance rate tuned to about $80\%$. At each temperature we generated several independent MC streams.
@@
We have generated the data sets listed in Tab. 1 and 2 according to the setup described above. These tables show the bare coupling in each case, the quark masses in terms of lattice spacing, and the number of configurations for each ensemble. Some ensembles (mainly those of $m_{l}=m_{s}/5$) were generated directly at some target temperatures, while others were generated over a wider range of temperatures for a given lattice spacing; for the latter, temperature interpolations are required for the continuum extrapolations of Sec. 3.3 to ensure the existence of at least three different lattice spacings in all cases. Consequently, we have highlighted in pink the ensembles from previous works of the HotQCD collaboration [21, 20] or of the TUMQCD collaboration [35], which are also used in the present analysis. Note that in the present work, as in the previous ones, the bare coupling is defined as $\beta=10/g_{0}^{2}$, which is a common notation for improved actions.
As in Refs. [21, 20], we use the following discretization of $E_{i}$ and $B_{i}$ fields:
$E_{i}(\mathbf{x},\tau)$ $=$ $U_{i}(\mathbf{x},\tau)U_{4}(\mathbf{x}+\hat{i},\tau)-U_{4}(\mathbf{x},\tau)U_{i}(\mathbf{x}+\hat{4})\,,$ (3.1)
@@
As briefly introduced in Sec. 2, direct measurements of the $G_{E,B}$ correlators are subject to substantial gauge fluctuations, resulting in a low signal-to-noise ratio. While the multi-level algorithm [29] successfully mitigates this problem in theories with strictly local interactions, such as quenched QCD, it is not applicable to real QCD with dynamical fermions. In addition, the magnetic correlator requires a renormalization method due to its non-trivial anomalous dimension [25]. To address both issues at once, we use the Symanzik-improved GF [45], also known as the "Zeuthen flow", which is a modification of the original proposals [31, 32, 33, 34] by incorporating an additional term that provides an $O(a^{2})$ improvement to the flow action.
@@
where $\nabla_{\mu}$ and $\nabla^{\star}_{\mu}$ stand for the usual lattice forward and backward covariant derivatives, respectively, and $S_{\rm{LW}}$ is the usual combination of plaquettes and rectangles of the Lüscher-Weisz action [37, 38]. Note that the flow time $\tau_{F}$ has units of length squared.
The main feature of Eq. (3.3) is to remove short-range fluctuations by smearing the gauge field over a region of radius $\sqrt{8\tau_{F}}$, which naturally becomes the inverse of a renormalization scale $\mu_{F}$. In essence, the Zeuthen Flow is, indeed, a renormalization scheme. For our analysis, we will focus on the region where $\sqrt{8\tau_{F}}/a\geq 1$, making sure that the flow time is always large enough to account for at least one lattice spacing. We also need to keep the flow scale smaller than the relevant physical scales of the calculations, and previous studies have found that an effective criterion is to use flow times such that $\sqrt{8\tau_{F}}/\tau\leq 0.30$ [16, 18, 46]. This is also consistent with the perturbative NLO calculations of the electric correlator, which predict a linear dependence in $\tau_{F}$ when the flow time is constrained to $\sqrt{8\tau_{F}}\leq\tau$ [47]. Small values of $\tau T$ are drastically affected by lattice artifacts, and our tests indicate that the amount of flow needed to eliminate these discretization effects is already close to the maximum flow scale $\sqrt{8\tau_{F}}/\tau=0.30$. Consequently, we will only consider the range $0.25\leq\tau T\leq 0.50$ for our extrapolations. Moreover, while we have set the minimum flow time to be at least of one lattice spacing, we have observed in practice that for $\tau T\geq 0.25$ the flow dependence of the correlators is rather small if the flow time is restricted to the range $\sqrt{8\tau_{F}}/\tau\geq 0.25$. Thus, our continuum/flow extrapolations will be performed within this specific range:
@@
The lattice generation and GF analysis presented in this work were performed using SIMULATeQCD [48, 49, 50] and the MILC code (with SciDAC libraries) at several high performance computing (HPC) centers around the world.

### v2

@@ 1 Introduction
The study of heavy quarks in the quark-gluon plasma (QGP) PhysRevLett.34.1353 ; STAR:2005gfr ; BRAHMS:2004adc is essential for understanding the general properties of strongly interacting matter produced in heavy-ion collisions (HIC). Heavy quarks, such as charm and bottom quarks, provide unique insights into the non-equilibrium dynamics of the QGP, as they are produced in the early stages of the collision and then propagate throughout the evolving plasma. Therefore, the study of interaction and equilibration of heavy quarks is crucial for understanding many properties of the QGP, such as its near-perfect fluid behavior Rapp:2018qla ; Dong:2019byy ; He:2022ywp .
@@
The interaction of heavy quarks with the medium can be effectively described by Langevin dynamics Moore:2004tg for sufficiently large masses, and their motion is characterized with two transport coefficients: the heavy-quark momentum diffusion coefficient $\kappa$ and the drag coefficient $\eta$. These transport coefficients play a central role, since they determine the momentum distributions of the heavy-flavor hadrons in the HIC, which have been precisely measured at RHIC PHENIX:2006iih ; STAR:2006btx and LHC ALICE:2012ab .
@@
The momentum diffusion coefficient $\kappa$ quantifies the average momentum transferred from the QGP to a heavy quark by random, uncorrelated in time, momentum kicks, with $3\kappa$ being the mean squared momentum transfer per unit time. Physically, it describes how much momentum a heavy quark loses and how fast it thermalizes in the plasma. For sufficiently heavy quarks, the relaxation time is expected to be $M/T$ times larger than that of light quarks Moore:2004tg ; PhysRevD.37.2484 , where $M$ is the mass of the heavy quark and $T$ is the temperature of the medium. Recent measurements have confirmed that this relaxation time is indeed shorter than the mean lifetime of the QGP (even with the enhancement of the prefactor $M/T$), suggesting a rather fast relaxation of heavy quarks and confirming the strongly interacting nature of the matter produced in HIC. Therefore, because of its relevance to the phenomenology of heavy ions, an accurate estimate of $\kappa$ has become a critical goal for both theoretical and experimental studies.
@@
Lattice field theory techniques provide a first-principles approach to the non-perturbative aspects of quantum chromodynamics (QCD), including the extraction of $\kappa$. However, the calculation is far from trivial, since it requires a spectral reconstruction of the correlation functions computed numerically on the lattice. Like all transport coefficients, the diffusion coefficient is directly encoded in the low-frequency behavior of the spectral function. It is precisely this inversion problem that makes the extraction extremely challenging, i.e., even when the correlators are determined precisely, the spectral functions extracted from them —and consequently the diffusion coefficient— can still carry large uncertainties.
@@
The heavy-quark diffusion coefficient has been calculated in perturbation theory PhysRevD.37.2484 ; Moore:2004tg ; Caron-Huot:2007rwy , but the calculation is reliable only for asymptotically large temperatures $T$. In the present work, we build on previous studies to extract the heavy-quark diffusion coefficient from lattice QCD Banerjee:2011ra ; Francis:2015daa ; Altenkort:2020fgs ; Banerjee:2022gen ; Brambilla:2022xbd ; Brambilla:2020siz ; Altenkort:2023oms ; Altenkort:2023eav ; Banerjee:2022uge . The HotQCD collaboration calculated $\kappa$ in a (2+1)-flavor lattice QCD framework for physical strange quarks and rather unphysical light quarks with mass $m_{l}=m_{s}/5$ Altenkort:2023eav ; Altenkort:2023oms corresponding to the pion mass of $m_{\pi}\simeq 320$ MeV, and for a temperature range between $T=195$ and $352$ MeV. The effect of light-quark masses is expected to be irrelevant for sufficiently large temperatures, but important for temperatures below $250$ MeV. Thus, the main goal of this work is to extend the previous calculation to an almost physical point, with light-quark masses of $m_{l}=m_{s}/20$ corresponding to a pion mass of $m_{\pi}\simeq 160$ MeV, and for a wider temperature range, from values close to the crossover temperature $T_{c}$ up to the GeV scale. The extension to high temperature is needed for making contact with perturbation theory.
@@ 2 Theoretical framework
Close to equilibrium, the response of strongly interacting matter to small perturbations can be studied in terms of a handful of transport coefficients, such as the charge diffusion coefficient, or the shear and bulk viscosities. These coefficients can be extracted from current-current correlation functions, where the type of current determines the specific coefficient to be computed. Among all the coefficients, the easiest one to extract is the heavy-quark momentum diffusion coefficient $\kappa$, which is formally defined as the mean-squared momentum transfer per unit time experienced by a heavy quark immersed in a QGP. Together with the equilibrium mean-squared momentum that the heavy quark carries in the medium, this determines how long it takes for the quark to thermalize in the QGP. For a more extended review see section 6 of Ding:2015ona .
@@
[…]
@@
where $\chi_{q}$ is the heavy quark number susceptibility, $q=c,b$ for charm and bottom, respectively. The determination of $D_{s}$ on the lattice through Eq. (2.3) is extremely challenging, as discussed in Ref. Petreczky:2008px . An alternative approach relies on the use of the heavy-quark effective theory (HQET) Eichten:1990vp as discussed in Refs. Caron-Huot:2009ncn ; Bouttefeux:2020ycy ; Casalderrey-Solana:2006fio . In the HQET approach, the current-current correlation function is expressed in terms of the chromo-electric and chromo-magnetic correlation functions Caron-Huot:2009ncn ; Bouttefeux:2020ycy :
@@
The $\omega\rightarrow 0$ limit of the corresponding spectral functions defines the transport coefficients $\kappa_{E}$ and $\kappa_{B}$ Caron-Huot:2009ncn ; Bouttefeux:2020ycy :
@@
Thus, the heavy-quark momentum diffusion coefficient, up to corrections ${\cal O}(T^{2}/M^{2})$, can be written as Bouttefeux:2020ycy :
@@
The chromo-electric and chromo-magnetic correlation functions are constructed in terms of gluonic operators, which suffer from strong short-distance fluctuations that yield low signal-to-noise ratio. Accurate estimation of the correlation functions (2.4) and (2.5) thus requires noise reduction techniques, such as multi-level updating Luscher:2001up . While the electric correlation function requires only a finite renormalization at finite lattice spacing, and is regulator independent within the $\overline{\rm MS}$} scheme Christensen:2016wdo , the magnetic correlator involves a non-trivial anomalous dimension Eichten:1990vp , and therefore it requires a specific renormalization method. On the lattice, we will use the gradient flow (GF) renormalization procedure Narayanan:2006rf ; Luscher:2009eq ; Luscher:2010iy ; Luscher:2011bx to solve both problems at once, which has been shown to be as good as multi-level algorithms for noise reduction in the quenched theory Altenkort:2020fgs ; Brambilla:2022xbd , while providing renormalized field operators for any flow time $\tau_{F}$ larger than the square of the lattice-spacing scale. We will explain our choice of renormalization method as well as a small description of our GF implementation in Sec. 3.
In both, the electric and the magnetic case, spectral reconstruction is a crucial and difficult aspect of the calculation. The main problem arises from the fact that the correlation functions are computed only at some discrete Euclidean time intervals $\tau$. Thus, to extract the spectral function $\rho(\omega,T)$ from these data, one has to use a theoretically motivated Ansatz. For this purpose we can use the perturbative QCD (pQCD) estimates of the spectral function. At LO and NLO Caron-Huot:2009ncn , the spectral function in pQCD is predicted to behave as $\rho(\omega,T)\propto\omega^{3}$ for large $\omega$, while the low-energy behavior of the spectral function directly yields the diffusion coefficient $\kappa$ as $\rho(\omega,T)\propto\kappa_{E,B}\omega$. To interpolate between these two regions, we will explore several methods, like for instance sharp, smooth, or polynomial transition between the IR and UV behavior of the spectral function. These methods are discussed in detail in Sec. 5. In order to determine $\kappa$ and $D_{s}$, we also need to determine the average thermal velocity, $\langle\mathbf{v}^{2}\rangle$} and averaged thermal momentum, $\langle\mathbf{p}^{2}\rangle$} of the heavy quark. Following previous works Altenkort:2023eav ; Altenkort:2023oms , in Sec. 5.3.2 we will determine these quantities in terms of a quasi-particle model (QPM).
@@ 3.1 Lattice Setup
$T$ [MeV] $\beta$ $a$ [fm] $a\crossproduct m_{s}$ $a\crossproduct m_{l}$ $N_{\sigma}$ $N_{\tau}$ # conf. streams
@@
$T$ [MeV] $\beta$ $a$ [fm] $a\crossproduct m_{s}$ $a\crossproduct m_{l}$ $N_{\sigma}$ $N_{\tau}$ # conf. streams
@@
Table 2: Input parameters of the lattice ensembles corresponding to light-quark masses $m_{l}=m_{s}/5$. We show in every case the temperature in MeV, the bare coupling with its corresponding lattice spacing in fm, the geometry $N_{\sigma}^{3}\crossproduct N_{\tau}$ and the number of configurations separated by 10 MC trajectories. The highlighted pink regions correspond to ensembles generated in previous works of the HotQCD collaboration Altenkort:2023oms ; Altenkort:2023eav or the TUMQCD collaboration Bazavov:2019qoo .
We have performed lattice calculations in (2+1)-flavor QCD with physical strange-quark mass $m_{s}$ over a temperature range from $T=149$ MeV to $10$ GeV using the Highly Improved Staggered Quarks (HISQ) action Follana:2006rc for the fermion sector and the tree-level improved Lüscher-Weisz action Luscher:1984xn ; Luscher:1985zq for the gauge fields. The calculations have been performed on $N_{\sigma}^{3}\crossproduct N_{\tau}$ boxes for two values of the light-quark mass: $m_{l}=m_{s}/20$, which corresponds to the pion mass of $160$ MeV in the continuum limit (and thus is close to the physical point), and $m_{l}=m_{s}/5$, corresponding to a pion mass of $320$ MeV. The lattice strange-quark mass was obtained from the parameterization of the line of constant physics (LCP) of HotQCD:2014kol . The lattice spacing, and thus the temperature scale, has been fixed through the $r_{1}$ scale determined in Refs. HotQCD:2014kol ; Bazavov:2017dsy with the value $r_{1}=0.3106(18)$ fm Bazavov:2010sb . Though a very recent determination of $r_{1}$ in physical units yielded a smaller central value Larsen:2025wvg , it is still compatible with the scale used in this paper.
The gauge configurations were generated using the usual Rational Hybrid Monte Carlo (RHMC) algorithm Clark:2003na ; Clark:2004cp , where each configuration is saved after 10 Monte Carlo (MC) trajectories with an acceptance rate tuned to about $80\%$. At each temperature we generated several independent MC streams.
@@
We have generated the data sets listed in Tab. 1 and 2 according to the setup described above. These tables show the bare coupling in each case, the quark masses in terms of lattice spacing, and the number of configurations for each ensemble. Some ensembles (mainly those of $m_{l}=m_{s}/5$) were generated directly at some target temperatures, while others were generated over a wider range of temperatures for a given lattice spacing; for the latter, temperature interpolations are required for the continuum extrapolations of Sec. 3.3 to ensure the existence of at least three different lattice spacings in all cases. Consequently, we have highlighted in pink the ensembles from previous works of the HotQCD collaboration Altenkort:2023eav ; Altenkort:2023oms or of the TUMQCD collaboration Bazavov:2019qoo , which are also used in the present analysis. Note that in the present work, as in the previous ones, the bare coupling is defined as $\beta=10/g_{0}^{2}$, which is a common notation for improved actions.
As in Refs. Altenkort:2023eav ; Altenkort:2023oms , we use the following discretization of $E_{i}$ and $B_{i}$ fields:
[…]
@@
As briefly introduced in Sec. 2, direct measurements of the $G_{E,B}$ correlators are subject to substantial gauge fluctuations, resulting in a low signal-to-noise ratio. While the multi-level algorithm Luscher:2001up successfully mitigates this problem in theories with strictly local interactions, such as quenched QCD, it is not applicable to real QCD with dynamical fermions. In addition, the magnetic correlator requires a renormalization method due to its non-trivial anomalous dimension Eichten:1990vp . To address both issues at once, we use the Symanzik-improved GF Ramos:2015baa , also known as the "Zeuthen flow", which is a modification of the original proposals Narayanan:2006rf ; Luscher:2009eq ; Luscher:2010iy ; Luscher:2011bx by incorporating an additional term that provides an $O(a^{2})$ improvement to the flow action.
@@
where $\nabla_{\mu}$ and $\nabla^{\star}_{\mu}$ stand for the usual lattice forward and backward covariant derivatives, respectively, and $S_{\rm{LW}}$ is the usual combination of plaquettes and rectangles of the Lüscher-Weisz action Luscher:1984xn ; Luscher:1985zq . Note that the flow time $\tau_{F}$ has units of length squared.
The main feature of Eq. (3.3) is to remove short-range fluctuations by smearing the gauge field over a region of radius $\sqrt{8\tau_{F}}$, which naturally becomes the inverse of a renormalization scale $\mu_{F}$. In essence, the Zeuthen Flow is, indeed, a renormalization scheme. For our analysis, we will focus on the region where $\sqrt{8\tau_{F}}/a\geq 1$, making sure that the flow time is always large enough to account for at least one lattice spacing. We also need to keep the flow scale smaller than the relevant physical scales of the calculations, and previous studies have found that an effective criterion is to use flow times such that $\sqrt{8\tau_{F}}/\tau\leq 0.30$ Altenkort:2020fgs ; Brambilla:2022xbd ; Eller:2018yje . This is also consistent with the perturbative NLO calculations of the electric correlator, which predict a linear dependence in $\tau_{F}$ when the flow time is constrained to $\sqrt{8\tau_{F}}\leq\tau$ Eller:2021qpp . Small values of $\tau T$ are drastically affected by lattice artifacts, and our tests indicate that the amount of flow needed to eliminate these discretization effects is already close to the maximum flow scale $\sqrt{8\tau_{F}}/\tau=0.30$. Consequently, we will only consider the range $0.25\leq\tau T\leq 0.50$ for our extrapolations. Moreover, while we have set the minimum flow time to be at least of one lattice spacing, we have observed in practice that for $\tau T\geq 0.25$ the flow dependence of the correlators is rather small if the flow time is restricted to the range $\sqrt{8\tau_{F}}/\tau\geq 0.25$. Thus, our continuum/flow extrapolations will be performed within this specific range:
@@
The lattice generation and GF analysis presented in this work were performed using SIMULATeQCD Mazur:2021zgi ; Altenkort:2021cvg ; HotQCD:2023ghu and the MILC code (with SciDAC libraries) at several high performance computing (HPC) centers around the world.